Solve the following problems, showing all of your calculations .

Assume the following:

GDP (Y) = 5,000

  • C (consumption)= 750 + 0.5(Y-T).
  • I (investment)= 1500 – 50r, where r is the real rate of interest in percent.
  • T (taxes)= 1,250
  • G(government spending) = 1,125
  1.    What are the equilibrium values of C, I and r?
  2. A technical innovation increases investment to the function I = 2,000 - 50r. What are the new equilibrium values of C, I, and r?





1
Expert's answer
2022-02-14T15:26:40-0500

"Y=C+I+G"

"C=750+0.5\u00d7(Y-T)"

"=750+0.5\u00d7(5000-1250)=2625"

"Y=2625+(1500-50r)+1125"

"5000=2625+(1500-50r)+1125"

"5000=5250-50r"

"50r=250"

"r=5"

"I=1500-50\u00d750=1250"

B)

"C=Y-I-G"

"C=5000-(2000-5r)-1125"

"C=1875+50r"

"C=1875+50\u00d750=2125"

"I=2000-50\u00d750=1750"